OpenStudy (anonymous):

Find functions f and g such that the given function h = f ∘ g. h(x) = (1 + 1/x)^x I would like a hint.

OpenStudy (anonymous):

Nooooooooo

OpenStudy (anonymous):

Wait a minute. If it is f(g(x)), then that should mean that every "x" of f(x) is turned into (1 + 1/x).

OpenStudy (anonymous):

Why isn't it that the x raised to the power is not turned into (1 + 1/x)?

OpenStudy (anonymous):

I was taught that wherever the x is, you have to replace it with the function g(x), if it is f(g(x)). In this case, I am confused because there is an "x" in the power that looks like it didn't get replaced.

myininaya (myininaya):

what?

OpenStudy (anonymous):

hello myininaya!

myininaya (myininaya):

hey

OpenStudy (anonymous):

lots of random answers this evening

OpenStudy (anonymous):

earlier too

myininaya (myininaya):

hey answer this one i'm having a brain fart

OpenStudy (anonymous):

how about \[g(x)=1+\frac{1}{x}\] \[h(x)=x^x\]

myininaya (myininaya):

we could do \[f(x)=(1+\frac{1}{x})^x, g(x)=x\] but i'm pretty sure this is not what he wants

OpenStudy (anonymous):

i like my answer more

myininaya (myininaya):

\[h(x)=x^x?\]

myininaya (myininaya):

do you mean f that won't work

myininaya (myininaya):

\[f(g(x))=f(1+\frac{1}{x})=(1+\frac{1}{x})^{1+\frac{1}{x}}\] not right

OpenStudy (anonymous):

\[g(x)=1+\frac{1}{x}\] \[f(x)=x^x\] \[f\circ g(x)=f(g(x))=f(1+\frac{1}{x})=(1+\frac{1}{x})^x\]

myininaya (myininaya):

that is actually what i said at first without thinking

myininaya (myininaya):

no you didn't replace x with g(x)

OpenStudy (anonymous):

i believe i did yes

myininaya (myininaya):

you didn't replace both

myininaya (myininaya):

you replaced one of the x's

myininaya (myininaya):

the base but not the exponent

OpenStudy (anonymous):

oooooooooooooooooooooh

OpenStudy (anonymous):

yes you are right!

myininaya (myininaya):

seriously what you have here though is exactly what i said and then he questioned me about it and i was like what and then i realized omg hes right

OpenStudy (anonymous):

so try \[f(x)=x^{\frac{1}{x-1}}\]

OpenStudy (anonymous):

bet it works now

OpenStudy (anonymous):

i just found the inverse of \[1+\frac{1}{x}\] and stuck that in the exponent

myininaya (myininaya):

nice!

OpenStudy (anonymous):

merci but nice of you to notice that i screwed up the first time!

myininaya (myininaya):

i gotta go eat :) gj satellite

OpenStudy (anonymous):

I'm confused.

myininaya (myininaya):

\[h=f(g(x))=f(1+\frac{1}{x})\] so we need to choose f such that we get \[f(1+\frac{1}{x})=(1+\frac{1}{x})^x\] if \[g(x)=1+\frac{1}{x}\] and we are suppose to plug in g(x) wherever we see an x in f(x), then we need \[h(x)=[g(x)]^{[v(x)]}\] we need to write v in terms of g we want \[v(x)=x=> g(x)=1+\frac{1}{v(x)}\] we need to solve this for v \[vg=v+1=> vg-v=1=>v(g-1)=1=> v=\frac{1}{g-1}\] so we have \[h(x)=f(g(x))=(g(x))^\frac{1}{g(x)-1}\] so this implies \[g(x)=1+\frac{1}{x}, f(x)=x^\frac{1}{x-1}\]